OpenStudy (anonymous):

If yz : zx : xy = 1 : 2 : 3, then (x/yz) / (y/zx) = ? If yz : zx : xy = 1 : 2 : 3, then (x/yz) / (y/zx) = ? @Mathematics

OpenStudy (anonymous):

\[(x/yz)\div (y /zx)\]

OpenStudy (anonymous):

Does anyone have any idea on How to do it..

OpenStudy (anonymous):

\[\frac{x}{yz}/\frac{y}{zx}=\frac{x}{yz}*\frac{zx}{y}=\frac{x^2z}{y^2z}=\frac{x^2}{y^2} \]

OpenStudy (anonymous):

\[\frac{yz}{zx}=\frac{1}{2}=\frac{y}{x}\] x=2y \[\frac{x^2}{y^2}=\frac{(2y)^2}{y^2}=\frac{2^2y^2}{y^2}=4\]

OpenStudy (anonymous):

Thanks Marina

OpenStudy (anonymous):

You are welcome. It was interesting question.

OpenStudy (anonymous):

I have another doubt. Would you mind answering it?

OpenStudy (anonymous):

Sorry, I have to go. Will be glad to help another time. Buy.

OpenStudy (anonymous):

Ok, bye